# Untestability of Average Slutsky Symmetry

### v1

@@ 4 Conclusion
In this paper, we addressed a question whether the symmetry of the average Slutsky matrix has empirical implications for cross-sectional demand distributions. We showed that the answer is negative: for any regular cross-sectional demand distributions, there exists a stochastic demand system that has a symmetric average Slutsky matrix. This result implies the untestability of individual Slutsky symmetry based on the average Slutsky matrix. Nonetheless, it does not rule out the possibility that individual Slutsky symmetry could have the empirical content through nonlinear relationship, which remains an important open question for future research.
@@ Appendix A Preliminaries
where is the outward unit normal vector on This PDE admits a unique solution such that by [nardi2015schauder]. Define a vector field as We first establish the regularity of
@@
is differentiable with a HΓΆlder continuous derivative uniformly over that is,
@@
in the weak sense (see Chapter 4 of [santambrogio2015optimal]), i.e, for the map is absolutely continuous and it holds
@@
Let Then, satisfies the continuity equation in the weak sense by the standard argument. By the uniqueness of the solution of the continuity equation (Theorem 4.4 of [santambrogio2015optimal]), we have Hence, for a fixed random variable the random process has a continuously differentiable sample path and
@@
(Theorem 3.9 of [teschl2012ordinary]). Let Letting
@@
By Gronwallβs inequality, we obtain
for some universal constant because is Lipschitz by Theorem 2.8 of [teschl2012ordinary] and is uniformly HΓΆlder by Lemma A.1. Hence, as which immediately implies is differentiable at with derivative
which is continuous in β
@@
Let and There exist functions and such that and hold.
@@
Proof.
By Theorem A.1, satisfies
where is the extension of by Lemma A.2. We first show that is smooth in the following sense.
Lemma A.3.
The vector field is differentiable in and and
1.
is continuous,
2.
is continuous,
3.
is continuous uniformly in and
4.
is continuous uniformly in
Consider the ODE
where
By Lemma A.3 and Theorem 3.12 of [teschl2012ordinary], this ODE admits a unique solution for each By Theorem 2.8 of [teschl2012ordinary], we have
for some constant Since the RHS converges to zero as by Lemma A.3, the map is continuous for each Finally, Lemma A.3 and Gronwallβs inequality show β
@@ Appendix C Omitted Proofs
For simplicity, we consider the case of There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Next, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Finally, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Let for and By construction, holds. The partial differentiability of in and follows by Theorem A.2. β
@@
Recall that the average Slutsky matrix at is
The symmetry of this matrix is equivalent to
where the RHS is identified from the marginal distributions because
holds. Since
holds, the symmetry is equivalent to
which completes the proof. β
@@
Let be a continuously differentiable function that has a compact support, is smooth in and
For given let
and
It is clear that in and on We also have
where the second equality holds by the integration by parts. Now, it is clear that satisfies the PDE (4). β
Proof of Lemma 3.4.
Since the second and third properties are obvious by construction, we shall prove the other two. The marginal compliance is shown using the continuity equation. Let be a test function. Then, the map is absolutely continuous and it holds
where the last equality holds by Lemma 3.3. For the partial differentiability, it suffices to show is differentiable, which is a consequence of Theorem A.2. Finally, the average symmetry condition (3) holds since
holds by Lemma 3.3 and the definition of β
Proof of Proposition 3.1.
@@
By the change-of-variable formula (see, for example, Lemma 2.62 of [sokolowski1992introduction]), solves
where and By (A.1.2), (A.1.6), and Theorem 2.1 of [kono2024well], is a unique solution satisfying and by his Theorem 3.1, an Schauder estimate
holds for some independent of 444Note that the coefficient is fixed in the original version of Theorem 3.1 of [kono2024well], but Theorem 6.30 of [gilbarg1977elliptic] ensures that depends on only through which is bounded by A.1.2. Using the equality we have
@@
for Since is continuous, so is by Bergeβs maximum theorem. Also, McShaneβs extension theorem implies that is Lipschitz uniformly over β
@@
The standard Schauder estimate shows the differentiability of by Assumption A.2.3. Since the derivative is continuous in and by Assumption A.2.3, the second and fourth bullet points are confirmed. β

### v2

@@ 4 Conclusion
In this paper, we have shown that the average Slutsky symmetry is not testable using cross-sectional demand data. To establish this, we explicitly derived the identified set of the average Slutsky matrix and demonstrated that it always contains a symmetric matrix. This finding implies that individual Slutsky symmetry cannot be tested through the average Slutsky matrix, although it does not rule out the possibility of testing the symmetry hypothesis without relying on the average. See [4] for recent developments. Furthermore, by imposing bounds on the income elasticity of demand, we showed that average Slutsky symmetry leads to a set of functional inequality constraints that are, in principle, testable. A promising direction for future research is to develop a statistical test for Slutsky symmetry based on these inequalities.
@@ Appendix A Preliminaries
where is the outward unit normal vector on This PDE admits a unique solution such that by [19]. Define a vector field as We first establish the regularity of
@@
is differentiable with a Hölder continuous derivative uniformly over that is,
@@
in the weak sense (see Chapter 4 of [20]), i.e, for the map is absolutely continuous and it holds
@@
Let Then, satisfies the continuity equation in the weak sense by the standard argument. By the uniqueness of the solution of the continuity equation (Theorem 4.4 of [20]), we have Hence, for a fixed random variable the random process has a continuously differentiable sample path and
@@
(Theorem 3.9 of [22]). Let Letting
@@
By Gronwall’s inequality, we obtain
for some universal constant because is Lipschitz by Theorem 2.8 of [22] and is uniformly Hölder by Lemma A.1. Hence, as which immediately implies is differentiable at with derivative
which is continuous in ∎
@@
Let and There exist functions and such that
hold.
@@ Appendix C Omitted Proofs
For simplicity, we consider the case of There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Next, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Finally, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Let for and By construction, holds. The partial differentiability of in and follows by Theorem A.2. ∎
@@
Let be an open ball such that in There exists a smooth nonnegative function such that and For each define a function as Then is and satisfies Moreover, the derivative is Lipschitz uniformly in due to the regularity of If we define then holds. Let where
We shall see that satisfies all the conditions of Lemma 3.2. By the property of and continuity of is Lipschitz uniformly in The third condition of (4) is satisfied because for
where the second equality holds by the integration by parts, and similarly, for
The second condition holds since vanishes on the boundary. The first condition holds since
@@
For the partial differentiability, it suffices to show is differentiable, which is a consequence of Theorem A.2. The marginal compliance (M) is shown using the continuity equation. Let be a test function. Then, the map is absolutely continuous and it holds
where the last equality holds by Lemma 3.2. By Lemma 3.2 and the definition of we have
which implies that condition (S) holds. ∎
Proof of Theorem 2.1.
It is easy to see the function is an element of Conversely, Lemma 3.3 shows that for a given there exists a -valued random demand function such that and for all For the second part of the theorem, if we define
then holds, and is symmetric for all ∎
@@
By the change-of-variable formula (see, for example, Lemma 2.62 of [21]), solves
where and By (A.1.2), (A.1.6), and Theorem 2.1 of [10], is the unique solution satisfying and by his Theorem 3.1, an Schauder estimate
holds for some independent of 333Note that the coefficient is fixed in the original version of Theorem 3.1 of [10], but Theorem 6.30 of [3] ensures that depends on only through which is bounded by A.1.2. Using the equality we have
@@
for Since is continuous, so is by Berge’s maximum theorem. Also, McShane’s extension theorem implies that is Lipschitz uniformly over ∎
@@
The standard Schauder estimate shows the differentiability of by Assumption A.2.3. Since the derivative is continuous in and by Assumption A.2.3, the second and fourth bullet points are confirmed. ∎